OpenStudy (anonymous):

determine the first five iterates of f(x)=1.5x(1-x) for x 0=0.8

OpenStudy (anonymous):

@GIL.ojei could you help me with this one as well?

OpenStudy (anonymous):

am sorry no

OpenStudy (anonymous):

It's okay. What about this one. Given f(x) = x^2+3 and g(x) = x+5/x fing (g o f)(4)

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yup.